OpenStudy (anonymous):

Find the distance between points M(6, 16) and Z(–1, 14) to the nearest tenth.

OpenStudy (anonymous):

Length of a line segment formula: AB √(X2-X1)² + (y2-y1)² MZ = √ (14-16)² + (-1-6)² MZ = √ (-2)² + (-7)² MZ = √ 4 + 49 MZ = √ 53 MZ = 7.280 Round to nearest tenth (i just googled it for you. i hope that's okay with you. i did what i could.)